OpenStudy (anonymous):

the coordinates of the vertex of the quadratic function f(x)=x^2-7x+12 are A. (7/2,-1/4) B. (-7/2,195/4) C. (7/2,-37/4) D. (7/2,-97/4) E. (6/7,330/49)

OpenStudy (saifoo.khan):

Do you know how to find the vertex? Hint: -b/2a ?

OpenStudy (anonymous):

-(-7)/2(1)=7/2 am i correct?

OpenStudy (saifoo.khan):

Yes. This gives us that.. .?

OpenStudy (anonymous):

its for x axis..how bout y axis?

hartnn (hartnn):

now plug in x=7/2 in f(x)=x^2-7x+12 to get your 'y' value :)

OpenStudy (anonymous):

yaa thank you i gt -1/4 @hartnn n @saifoo.khan

hartnn (hartnn):

thats correct :)